I have to solve the equation x^2+ 12x-64=0 using the 6 step Indian process and I am not coming up with good numbers
Found 2 solutions by mananth, Edwin McCravy:
Answer by mananth(16946)

x^2+ 12x-64=0
x^2+16x-4x-64=0
x(x+16)-4(x+16)=0
(x+16)(x-4)=0
x=-16 OR x=4

Answer by Edwin McCravy(20055)


The other tutor did the problem correctly but not by
the 6-step Indian process. Here it is done by the 
6-step Indian process.

(1) Get the constant term off the left side of
    the equation. 

    We add 64 to both sides:

    x² + 12x - 64 = 0
         x² + 12x = 64 


(2) Multiply each term in the equation by four 
    times the coefficient of the x² term. 

         x² + 12x = 64

    The coefficient of x² is 1
    Four times 1 is 4
    We multiply each term by 4

        4x² + 48x = 256      
     

(3) Square the coefficient of the original x term 
    and add it to both sides of the equation. 

    The coefficient of the original x term was 12 
    Square 12, get 144.
    Add 144 to both sides:

          4x² + 48x = 256
    4x² + 48x + 144 = 256 + 144
    4x² + 48x + 144 = 400
    The left side is a perfect square and may be
    written as the square root of the the first
    term plus the square root of the third term
    in parentheses squared: 

         (2x + 12)² = 400

(4) Take the square root of both sides. 

            2x + 12 = ±20

(5) Set the left side of the equation equal to the 
    positive square root of the number on the right 
    side and solve for x. 

            2x + 12 = 20
                 2x = 8
                  x = 4    


(6) Set the left side of the equation equal to the 
    positive square root of the number on the right 
    side and solve for x. 

            2x + 12 = -20
                 2x = -32
                  x = -16


Edwin